# PersuasiveToM: A Benchmark for Evaluating Machine Theory of Mind in Persuasive Dialogues

### v1

@@ Abstract
The ability to understand and predict the mental states of oneself and others, known as the Theory of Mind (ToM), is crucial for effective social interactions. Recent research has emerged to evaluate whether Large Language Models (LLMs) exhibit a form of ToM. Although recent studies have evaluated ToM in LLMs, existing benchmarks focus predominantly on physical perception with principles guided by the Sally-Anne test in synthetic stories and conversations, failing to capture the complex psychological activities of mental states in real-life social interactions. To mitigate this gap, we propose PersuasiveToM, a benchmark designed to evaluate the ToM abilities of LLMs in persuasive dialogues. Our framework introduces two categories of questions: (1) ToM Reasoning, assessing the capacity of LLMs to track evolving mental states (e.g., desire shifts in persuadees), and (2) ToM Application, evaluating whether LLMs can take advantage of inferred mental states to select effective persuasion strategies (e.g., emphasize rarity) and evaluate the effectiveness of persuasion strategies. Experiments across eight state-of-the-art LLMs reveal that while models excel on multiple questions, they struggle to answer questions that need tracking the dynamics and shifts of mental states and understanding the mental states in the whole dialogue comprehensively. Our aim with PersuasiveToM is to allow an effective evaluation of the ToM reasoning ability of LLMs with more focus on complex psychological activities. Our code is available at https://github.com/Yu-Fangxu/PersuasiveToM.
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) have demonstrated performance comparable to humans in problem-solving tasks. To assess whether LLMs exhibit high-level reasoning abilities regarding mental states, various studies have proposed benchmarks to evaluate their capacity to handle ToM tasks. A foundational concept in these benchmarks is the Sally-Anne test Baron-Cohen et al. (1985), which has inspired the development of ToM evaluation frameworks Gu et al. (2024); He et al. (2023). In this test, Anne secretly moves an object initially known to both Sally and Anne, leading Sally to hold a false belief about the object’s location. The task requires participants to reason about "Where will Sally look for the object?". Although this test assesses basic awareness, it falls short in capturing the complex dynamics of mental states in real-life social interactions and may not fully reflect ToM abilities in practical scenarios. To better simulate real-world social contexts, some benchmarks have been developed around communication scenarios Kim et al. (2023); Chan et al. (2024). However, these benchmarks focus primarily on inferring the scope of information awareness and often involve characters in equal positions engaged in information exchange. This limits the ToM evaluation of reasoning psychological states. Additionally, current ToM benchmarks often overlook the critical step of applying ToM reasoning to predict actions, which is a key component of advanced social cognition.
To address these limitations, we introduce PersuasiveToM, a benchmark designed to evaluate LLMs in real-life social scenarios, focusing on ToM reasoning and its practical applications. PersuasiveToM is built around persuasive dialogue, placing LLMs in a realistic social interaction scenario characterized by asymmetrical social status for complex psychological activities. Unlike traditional benchmarks that focus on information awareness (e.g., the location of an object), PersuasiveToM draws inspiration from the Belief-Desire-Intention (BDI) model Bratman (1987); Georgeff et al. (1999) to shift the focus from physical perception to psychological states, such as a character’s attitude toward an event, which are more complex to reason about. Furthermore, PersuasiveToM goes beyond reasoning about mental states by assessing how well LLMs can predict actions (e.g., persuasion strategies) based on their understanding of mental states and evaluate the effectiveness of these strategies based on the persuadee’s reactions.
@@ 2 Related Works
Existing ToM evaluation benchmarks for LLMs are mainly text story-based QA forms Gandhi et al. (2024); Le et al. (2019); Kim et al. (2023); He et al. (2023); Gu et al. (2024) with multi-modal extensions Jin et al. (2024a); Shi et al. (2024), which adapt or extend the Sally-Anne test Baron-Cohen et al. (1985). The questions in these benchmarks ask a model to select the true hypothesis of a person’s knowledge and belief based on a given premise. However, such benchmarks are inherently suffer from shortcuts and spurious correlations Sclar et al. (2023); Ullman (2023); Shapira et al. (2023); Ma et al. (2023) Moreover, most story-based benchmarks simply ask characters where the objects are located, which lacks a real social interaction scenario. To further address this weakness, Hou et al. (2024) evaluates the ToM of agents in a situated environment, Chan et al. (2024) evaluates the ability of ToM in a negotiation environment, yet it is limited to bargaining specific resources, water, food, and firewood. Persuasion scenarios involve more complex psychological activities and unequal character relationships that lead to differences in the mental states of both parties. Therefore, this work introduces PersuasiveToM, which assesses an LLM’s ability to accurately reason the mental states of individuals in persuasive multi-turn dialogues. We also evaluate whether they can appropriately apply the understanding of mental states for persuasion strategy prediction and evaluation, which connects ToM reasoning with decision-making in social scenarios.
@@
Persuasive dialogues aim to influence the beliefs, attitudes, or behaviors of individuals through communication strategies Shi et al. (2020). Recent works have tried to develop datasets or facilitate LLMs with persuasion techniques to achieve specific goals. Previous datasets are constructed by crowd-sourcing Wang et al. (2019) or synthesizing with LLMs Zhou et al. (2023); Jin et al. (2024b). Many of the previous works build an effective persuasive dialogue system from emotional influence Samad et al. (2022), social facts Chen et al. (2022), and strategies Tian et al. (2020); Jin et al. (2023). Previous persuasion techniques have been widely adopted to jailbreak LLMs Zeng et al. (2024), mislead LLMs Xu et al. (2023), as well as for information retrieval Furumai et al. (2024). A similar work Sakurai and Miyao (2024) evaluates the intention detection abilities of LLMs in persuasive dialogues; however, PersuasiveToM introduces a more comprehensive benchmark to assess the ToM abilities of LLMs in such contexts.
@@ 3.1 Overview
In constructing the PersuasiveToM benchmark, we aim to evaluate the Theory of Mind (ToM) abilities of LLMs in dynamic, multi-turn persuasive dialogues with asymmetric social status, which leads to different mental states. Our dataset construction focuses on two primary dimensions: ToM Reasoning (§3.2) and ToM Application (§3.3). ToM Reasoning assesses the models’ ability to track and understand the evolving mental states of both the persuader and the persuadee, including desires, beliefs, and intentions. ToM Application evaluates whether LLMs can leverage their inferred understanding of mental states to select and apply effective persuasion strategies, such as predicting the next strategy or judging the effectiveness of a given strategy based on the persuadee’s reactions. There are several key considerations when constructing PersuasiveToM. (1) The dataset should contain diverse persuasive domains (e.g., life, education, technology, etc.) to ensure a comprehensive evaluation in the social context. (2) The mental states should be changed after multi-turn interactions to assess whether LLMs can track the shift in the dialogue. (3) The mental states of both parties should be asymmetric (e.g., persuaders has relatively stable mental states with the guidance of static desire for the goal, yet persuadee’s mental states shift drastically under the proactive persuasion by persuaders.)
Table 1 presents an example of PersuasiveToM, illustrating how mental states are tracked and how ToM reasoning is applied across multiple turns in a persuasive dialogue. The example demonstrates the dynamic nature of desire shifts, belief updates, and intention inferences, as well as the application of these mental states to predict and evaluate persuasion strategies. This example highlights the complex psychological activities of real-world persuasive interactions and the challenges LLMs face in accurately reasoning in such social contexts.
@@ 3.2 ToM Reasoning
Desire represents a motivational state that drives behavior but does not necessarily imply a firm commitment Malle and Knobe (2001); Kavanagh et al. (2005). Desires are seen as either fulfilled or unfulfilled which is different form beliefs that are evaluated in terms of truth or falsity. In PersuasiveToM, we evaluate LLMs’ ability to comprehend and track the evolution of desires in both persuaders and persuadees. For the persuader, the desire is typically static, representing their goal (e.g., Persuade Alice to join the botanical garden tour). For the persuadee, however, desires are dynamic and shift in response to the persuader’s tactics (e.g., Alice’s initial desire to shop transforms into a willingness to compromise). To assess this, we design Desire Questions that probe two key aspects: (1) Can LLMs consistently identify the persuader’s static desire throughout the dialogue? (2) Can LLMs track the dynamics of the persuadee’s desire shifting from refusal or disinterest to being persuaded?
Belief Reasoning.
Belief is a cognitive state where an individual holds a particular perspective, attitude, or viewpoint regarding a given proposition or idea. In PersuasiveToM, beliefs refer to understanding and reasoning the attitudes of the opponent toward the goal, which is explicitly or implicitly expressed in the dialogue. For example, in Turn 1, Bob believes Alice is hesitant about the tour, while Alice believes Bob is enthusiastic. By Turn 3, Bob’s belief shifts to thinking Alice is considering the idea, while Alice becomes more informed about the garden’s history. Belief Questions ask LLMs to infer what the will <persuader/persuadee> believe <persuadee/persuader>’s attitude towards the persuasion goal. These questions require models to understand cues in utterances and update beliefs dynamically as the dialogue progresses.
@@
Table 2: Intention mapping from the persuasive principles. Refer to Appendix C for definitions of persuasive principles.
@@
Intentions represent deliberate commitments to pursue specific goals based on desires and beliefs, often linked to tangible actions aimed at achieving those objectives. Intentions have been extensively studied in psychology tests such as action prediction Phillips et al. (2002) and behavioral re-enactment Meltzoff (1995). Inspired by persuasion principles Cialdini and Goldstein (2004); Cialdini and Cialdini (2007), we develop a mapping from persuasion principles to intentions, as shown in table 2. In persuasive dialogue, persuasive strategies have a strong association with intentions Wang et al. (2019). In PersuasiveToM, we collect the persuasive strategies from the PersuasionDaily dataset and their corresponding utterances for prompting the LLMs to choose the most appropriate intentions from table 2. The details of the extraction is recorded in Appendix A. For the persuader, we ask LLMs to choose the most appropriate intention from the six designed intention choices. For the persuadee, intentions are extracted by LLMs from utterances that are unrelated to persuasion intention.
@@ 3.3 ToM Application
This question involves asking which persuasion strategy the persuader is likely to employ next from a set of possible strategies. To answer these questions correctly, LLMs need to reason over the dialogue to infer the mental states of characters and predict what is the likely next prediction strategy to further influence the persuadee’s beliefs, desires, and intentions, ultimately achieving the desired persuasion outcome.
@@
The judgement question specifies the correct strategy was taken, and asks LLMs if the selected strategy is effective for persuasion. Answering such questions requires reasoning about the beliefs and intentions of the persuadee. Only by accurately inferring the persuadee’s mental state can properly determine whether the persuasion strategies should be employed to convince the persuadee.
3.4 Statistics and Analysis
Data Source.
PersuasiveToM is annotated on the multi-turn persuasive dialogue dataset DailyPersuasion Jin et al. (2024b). Each instance in DailyPersuasion is an $N$-round alternating dialogue $D=[(u_{1}^{a},u_{1}^{b},s_{1}^{a}),(u_{2}^{a},u_{2}^{b},s_{2}^{a}),...,(u_{N}^{a},u_{N}^{b},s_{N}^{a})]$ between the persuader $a$ and the perusadee $b$, and accompanied with a persuasion strategy $s_{i}^{a}$. persuadee $b$ has different desire from $a$ initially, after multi-turn persuasion, persuadee $b$ changes the mind to agree or consider the proposal of $a$.
Statistics.
In Table 3, we present the data statistics of PersuasiveToM. As shown in Figure 2, PersuasiveToM includes diverse domains. These real-life domains are crucial for comprehensively evaluating LLMs in social interactions. We sample 15 dialogues from each domain to form the dataset in PersuasiveToM. We create multi-choice questions by either prompting GPT-4o to generate semantically different choices or randomly selecting three distractors from the predefined pool. Refer to Appendix A for more details.
@@ 3.4 Statistics and Analysis
Model Desire Belief Intention Desire Belief Intention State Pred Judgement
@@
Human 100.00 92.31 78.02 84.62 87.91 94.50 86.81 97.83
LLaMa-3.1-8B-Instruct 58.78 66.28 40.85 69.91 71.82 86.77 62.22 96.37
Qwen2.5-7B-Chat 96.33 82.37 45.64 65.15 79.06 85.84 58.94 82.99
Gemma-2-9b-it 98.20 82.52 45.98 61.37 64.07 82.31 65.02 56.93
GLM4-9B-Chat 89.45 73.64 41.24 65.23 66.82 85.35 61.29 91.47
Mixtral-8x7B-Instruct 92.69 67.56 42.95 69.74 72.56 85.10 61.15 96.81
InternLM-2.5-7B-Chat 83.80 69.65 39.87 71.45 69.11 87.31 65.07 89.41
GPT-4o-mini 82.87 69.98 45.66 70.72 76.47 87.56 65.50 92.65
GPT-4o 98.75 89.49 46.02 50.10 80.03 88.78 73.00 97.55
LLaMa-3.1-8B-Instruct + CoT 59.75 69.81 43.11 68.64 73.45 85.35 61.98 77.27
Qwen2.5-7B-Chat + CoT 74.57 80.30 46.07 67.63 79.40 84.62 66.39 96.22
Gemma-2-9b-it + CoT 95.91 83.30 45.30 64.98 66.55 81.67 67.11 66.55
GLM4-9B-Chat + CoT 87.46 69.53 45.56 59.94 71.07 83.93 63.98 91.42
Mixtral-8x7B-Instruct + CoT 92.64 72.05 45.05 67.81 71.28 85.84 65.90 92.65
InternLM-2.5-7B-Chat + CoT 93.57 69.92 39.41 50.79 65.34 84.32 59.78 77.90
GPT-4o-mini + CoT 93.42 71.82 45.55 66.29 78.04 85.55 66.08 89.61
GPT-4o + CoT 96.51 83.57 46.10 68.72 83.56 87.54 77.06 95.88
Table 4: Results of LLMs on PersuasiveToM. Bold font and underlining indicate the best and second-best performance respectively.

### v2

@@ Abstract
The ability to understand and predict the mental states of oneself and others, known as the Theory of Mind (ToM), is crucial for effective social scenarios. Although recent studies have evaluated ToM in Large Language Models (LLMs), existing benchmarks focus on simplified settings (e.g., Sally-Anne-style tasks) and overlook the complexity of real-world social interactions. To mitigate this gap, we propose PersuasiveToM, a benchmark designed to evaluate the ToM abilities of LLMs in persuasive dialogues. Our framework contains two core tasks: ToM Reasoning, which tests tracking of evolving desires, beliefs, and intentions; and ToM Application, which assesses the use of inferred mental states to predict and evaluate persuasion strategies. Experiments across eight leading LLMs reveal that while models excel on multiple questions, they struggle with the tasks that need tracking the dynamics and shifts of mental states and understanding the mental states in the whole dialogue comprehensively. Our aim with PersuasiveToM is to allow an effective evaluation of the ToM reasoning ability of LLMs with more focus on complex psychological activities. Our code is available at https://github.com/Yu-Fangxu/PersuasiveToM.
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) have demonstrated performance comparable to humans in problem-solving tasks. To assess whether LLMs exhibit high-level reasoning abilities regarding mental states, various studies have proposed benchmarks to evaluate their capacity to handle ToM tasks. A foundational concept in these benchmarks is the Sally-Anne test Baron-Cohen et al. (1985), which has inspired the development of ToM evaluation frameworks Gu et al. (2024); He et al. (2023). In this test, Anne secretly moves an object initially known to both Sally and Anne, leading Sally to hold a false belief about the object’s location. The task requires participants to answer "Where will Sally look for the object?". Although this test assesses perception of the physical world, it fails to capture the complex dynamics of mental states in real-life social interactions and may not fully reflect ToM abilities in practical scenarios. To better simulate real-world social contexts, some benchmarks have been developed around communication scenarios Kim et al. (2023); Chan et al. (2024). However, these benchmarks still focus primarily on inferring the scope of information regarding the physical world. This limits the ToM evaluation of understanding psychological states. Furthermore, current ToM benchmarks often overlook the critical step of applying ToM reasoning to predict actions, which is a key component of advanced social cognition.
To address these limitations, we introduce PersuasiveToM, a benchmark designed to evaluate LLMs’ Theory of Mind capabilities specifically in realistic social interactions. Unlike previous benchmarks focused on inferring information about the physical world (e.g., object locations in Sally-Anne tests), PersuasiveToM centers on understanding complex psychological states, such as a character’s attitude towards an event. Inspired by the Belief-Desire-Intention (BDI) model Bratman (1987); Georgeff et al. (1999), PersuasiveToM uses persuasive dialogue scenarios, characterized by asymmetric social status, to generate different psychological states for both parties. In addition, beyond assessing ToM reasoning, PersuasiveToM evaluates the application of this understanding: assessing how well LLMs can predict actions (e.g., persuasive strategies) based on inferred mental states, and evaluating the effectiveness of persuasive strategies based on the persuadee’s reactions.
@@
[…]
ToMi
Hi-ToM
ToMBench
OpenToM
Big-ToM
FANToM
NegotiationToM
PersuasiveToM
Table 1: Comparison of PersuasiveToM with existing ToM datasets. In the header, Psychological ToM refers to testing ToM abilities in characters’ mental states of the psychological activities.
@@ 2 Related Works
Existing ToM evaluation benchmarks for LLMs are mostly text story-based QA forms Gandhi et al. (2024); Le et al. (2019); Kim et al. (2023); He et al. (2023); Gu et al. (2024), with some multi-modal extensions Jin et al. (2024a); Shi et al. (2024), which adapt the Sally-Anne test Baron-Cohen et al. (1985). These benchmarks ask models to determine the true beliefs or knowledge of individuals based on a given premise Sclar et al. (2023); Ullman (2023); Shapira et al. (2023); Ma et al. (2023). Story-based benchmarks focus on reasoning about mental states of the physical world without on psychological states, and applying ToM for decision-making in real-world social interactions. To address this, Hou et al. (2024) evaluates ToM in situated environments, and Chan et al. (2024) focuses on negotiation scenarios. However, the latter is limited to bargaining specific resources, while ToMATO Shinoda et al. (2025) uses persuasive dialogues but focuses solely on donation requests, limiting its diversity. For decision-making, BigToM Gandhi et al. (2024) assesses action prediction in narratives, but without grounding in interactive social interactions. Persuasion involves complex psychological dynamics and asymmetric relationships, providing a rich testbed for assessing ToM in social interactions. Built upon persuasive dialogues, PersuasiveToM is designed to evaluate an LLM’s ability to reason about the mental states of individuals and apply this understanding to predict and assess persuasion strategies, bridging ToM reasoning with decision-making in social interactions. Table 1 compares PersuasiveToM with existing ToM benchmarks.
@@
Persuasive dialogues aim to influence the beliefs, attitudes, or behaviors of individuals through communication strategies Shi et al. (2020). Recent works have tried to develop datasets or facilitate LLMs with persuasion techniques to achieve specific goals. Previous datasets are constructed by crowd-sourcing Wang et al. (2019) or synthesizing with LLMs Zhou et al. (2023); Jin et al. (2024b). Many of the previous works build an effective persuasive dialogue system from emotional influence Samad et al. (2022), social facts Chen et al. (2022), and strategies Tian et al. (2020); Jin et al. (2023). Previous persuasion techniques have been widely adopted to jailbreak LLMs Zeng et al. (2024), mislead LLMs Xu et al. (2023), as well as for information retrieval Furumai et al. (2024). A similar work Sakurai and Miyao (2024) evaluates the intention detection abilities of LLMs in persuasive dialogues; however, PersuasiveToM introduces a more comprehensive benchmark to assess the ToM abilities of LLMs in such contexts. In addition, we discuss the practical connection with dialogue generation and personalization in Appendix A.
@@ 3.1 Overview
We construct PersuasiveToM benchmark to evaluate the Theory of Mind (ToM) abilities of LLMs in dynamic, multi-turn persuasive dialogues with asymmetric social status, which give rise to distinct mental states. PersuasiveToM focuses on two core dimensions: ToM Reasoning3.3), assessing whether models can track evolving desires, beliefs, and intentions of both persuader and persuadee; and ToM Application3.4), evaluating whether LLMs can use these inferred states to predict or assess persuasion strategies. Kye design considerations include: (1) coverage of diverse domains (e.g., life, education, technology, etc.) to ensure a comprehensive evaluation in the social context. (2) evolving mental states across multi-turn interactions to assess whether LLMs can track the shift in the dialogue. (3) asymmetric roles have different mental states (e.g., stable goals for persuaders vs. shifting states for persuadees under active persuasion).
Table 2 shows a PersuasiveToM example, illustrating how mental states—desires, beliefs, and intentions—are tracked and inferred across turns, and how they inform the prediction and evaluation of persuasion strategies. This highlights the dynamic nature of real-world persuasion and the reasoning challenges it poses for LLMs.
3.2 Data Source
PersuasiveToM is annotated on the multi-turn persuasive dialogue dataset DailyPersuasion Jin et al. (2024b). Each instance in DailyPersuasion is an $N$-round alternating dialogue $D=[(u_{1}^{a},u_{1}^{b},s_{1}^{a}),(u_{2}^{a},u_{2}^{b},s_{2}^{a}),...,(u_{N}^{a},u_{N}^{b},s_{N}^{a})]$ between the persuader $a$ and the perusadee $b$, and accompanied with a persuasion strategy $s_{i}^{a}$. Persuadee $b$ has a different desire from $a$ initially, after multi-turn persuasion, persuadee $b$ changes the mind to agree or consider the proposal of $a$.
@@ 3.3 ToM Reasoning
Desire represents a motivational state that drives behavior but does not necessarily imply a firm commitment Malle and Knobe (2001); Kavanagh et al. (2005). Desires are seen as either fulfilled or unfulfilled which is different form beliefs that are evaluated in terms of truth or falsity. In PersuasiveToM, we evaluate LLMs’ ability to comprehend and track the evolution of desires in both persuaders and persuadees. For the persuader, the desire is typically static, representing their goal (e.g., persuade Alice to join the botanical garden tour). For the persuadee, however, desires are dynamic and shift in response to the persuader’s tactics (e.g., Alice’s initial desire to shop transforms into a willingness to compromise). To assess this, we design Desire Questions that probe two key aspects: (1) Can LLMs consistently identify the persuader’s static desire throughout the dialogue? (2) Can LLMs track the dynamics of the persuadee’s desire shifting from refusal or disinterest to being persuaded? For evaluation, we annotate the persuader’s desire questions as the persuasive goal in DailyPersuasion and use LLMs to annotate whether the persuadee is ultimately persuaded. See Appendix B for details.
@@
Table 3: Intention mapping from the persuasive principles. Refer to Appendix D for definitions of persuasive principles.
Belief Reasoning.
Belief is a cognitive state where an individual holds a particular perspective, attitude, or viewpoint regarding a given proposition or idea. In PersuasiveToM, beliefs refer to understanding and reasoning the attitudes of the opponent toward the goal, which is explicitly or implicitly expressed in the dialogue. For example, in Turn 1, Bob believes Alice is hesitant about the tour, while Alice believes Bob is enthusiastic. By Turn 3, Bob’s belief shifts to thinking Alice is considering the idea, while Alice becomes more informed about the garden’s history. Belief Questions ask LLMs to infer what will <persuader/persuadee> believe <persuadee/persuader>’s attitude towards the persuasion goal. These questions require models to understand cues in utterances and update beliefs dynamically as the dialogue progresses. We annotate the attitudes as the tone of each utterance of both persuaders and persuadees in DailyPersuasion.
@@
Intentions represent deliberate commitments to pursue specific goals based on desires and beliefs, often linked to tangible actions aimed at achieving those objectives. Intentions have been extensively studied in psychology tests such as action prediction Phillips et al. (2002) and behavioral re-enactment Meltzoff (1995). Inspired by persuasion principles Cialdini and Goldstein (2004); Cialdini and Cialdini (2007), we develop a mapping from persuasion principles to intentions, as shown in Table 3. In persuasive dialogue, persuasive strategies have a strong association with intentions Wang et al. (2019). In PersuasiveToM, we collect the persuasive strategies from the DailyPersuasion dataset and their corresponding utterances for prompting the LLMs to choose the most appropriate intentions from table 3. The details of the extraction are recorded in Appendix B. For the persuader, we ask LLMs to choose the most appropriate intention from the six designed intention choices. For the persuadees, intentions are summarized and extracted by LLMs from their utterances.
@@ 3.4 ToM Application
This question involves asking which persuasion strategy the persuader is likely to employ next from a set of possible strategies. To answer these questions correctly, LLMs need to reason over the dialogue to infer the mental states of characters and predict what the likely next prediction strategy is to further influence the persuadee’s beliefs, desires, and intentions, ultimately achieving the desired persuasion outcome.
@@
The judgment question specifies that the correct strategy was taken, and asks LLMs if the selected strategy is effective for persuasion. Answering such questions requires reasoning about the beliefs and intentions of the persuadee. Only by accurately inferring the persuadee’s mental state can one properly determine whether the persuasion strategies should be employed to convince the persuadee.
3.5 Statistics
In Table 4, we present the data statistics of PersuasiveToM. As shown in Figure 2, PersuasiveToM includes diverse domains. These real-life domains are crucial for comprehensively evaluating LLMs in social interactions. We sample 15 dialogues from each domain to form the dataset in PersuasiveToM. We create multi-choice questions by either prompting GPT-4o Hassany et al. (2025) to generate semantically different choices or randomly selecting three distractors from the predefined pool. Refer to Appendix B for more details.
@@ 3.5 Statistics
Model Desire Belief Intention Desire Belief Intention Strategy Pred Judgement
@@
Human 100.00 92.31 78.13 84.62 87.85 94.42 86.80 97.95
LLaMa-3.1-8B-Instruct 59.78±1.41 65.84±0.63 41.45±0.85 68.09±2.57 71.54±0.40 83.67±4.38 61.69±0.76 93.58±3.95
Qwen2.5-7B-Instruct 91.08±7.42 83.09±1.01 46.09±0.63 64.76±0.55 79.00±0.08 82.77±4.35 63.23±6.06 95.63±0.28
Gemma-2-9b-it 96.06±3.02 83.36±1.20 45.42±0.78 62.87±2.12 64.62±0.78 80.24±2.93 64.75±0.39 63.23±8.92
GLM4-9B-Chat 87.84±2.28 69.00±6.55 40.82±0.59 63.18±2.90 67.25±0.61 83.54±2.56 61.10±0.27 93.38±2.70
Mixtral-8x7B-Instruct 93.10±0.59 71.03±4.90 41.68±1.80 68.27±2.08 70.09±3.50 83.92±1.67 63.06±2.70 95.36±2.05
InternLM-2.5-7B-Chat 83.73±0.09 70.46±1.14 39.77±0.13 71.94±0.69 69.50±0.54 85.29±2.86 64.01±1.50 74.29±0.61
GPT-4o-mini 94.70±0.22 70.66±0.96 45.19±0.66 70.78±0.09 78.25±2.52 86.62±1.33 66.13±0.90 91.52±1.60
GPT-4o 95.56±4.51 89.47±0.02 46.28±0.37 67.66±2.23 81.73±2.41 87.81±1.38 73.55±0.78 96.38±1.66
LLaMa-3.1-8B-Instruct + CoT 59.83±0.11 68.91±1.27 41.47±2.32 66.50±3.02 70.53±4.13 84.30±1.48 62.97±1.40 78.99±2.43
Qwen2.5-7B-Instruct + CoT 86.64±2.30 82.66±3.34 45.69±0.54 66.61±1.44 78.77±0.89 83.41±1.72 65.94±0.64 94.56±2.35
Gemma-2-9b-it + CoT 95.72±0.26 83.35±0.07 44.89±0.58 63.46±2.15 66.07±0.68 78.40±4.62 67.56±0.64 67.28±1.04
GLM4-9B-Chat + CoT 86.58±1.24 67.01±3.56 44.49±1.51 59.72±0.30 69.34±2.45 82.06±2.64 63.21±1.09 92.91±2.11
Mixtral-8x7B-Instruct + CoT 92.47±0.23 73.88±2.59 42.15±4.10 66.09±2.43 70.64±0.91 84.40±2.04 66.71±1.14 92.01±0.91
InternLM-2.5-7B-Chat + CoT 89.50±5.76 69.25±0.95 39.34±0.09 49.72±1.52 64.48±1.21 80.48±5.42 56.98±3.97 76.72±1.66
GPT-4o-mini + CoT 91.65±2.51 71.33±0.69 46.15±0.85 68.56±3.22 76.27±2.51 83.63±2.71 67.16±1.52 89.22±0.55
GPT-4o + CoT 93.80±3.83 85.50±2.74 46.92±1.17 67.56±1.64 84.90±1.89 87.17±0.52 76.90±0.23 95.53±0.49
Table 5: Results of LLMs on PersuasiveToM. Bold font and underlining indicate the best and second-best performance, respectively.